OpenStudy (anonymous):

A bacteria culture initially contains 2000 bacteria and doubles every half hour. a) Find an expression for the number of bacteria after hours.

OpenStudy (anonymous):

@abb0t You trolling? lol

OpenStudy (abb0t):

OMG. LOL. i should be in bed. My final is in like 5 hrs!!!!

OpenStudy (anonymous):

Standard equation for an exponential function base 2 (aka it doubles) \[ F(t) = C\cdot 2^{kt} \] "A bacteria culture initially contains 2000" Meaning:\[ F(0) = C\cdot 2^{k(0)} = 2000 \] "doubles every half hour" Meaning:\[ F(1/2) = C\cdot 2^{k(1/2)} = 2\cdot 2000 = 4000 \]Two equations and two variables \(C,k\)

OpenStudy (anonymous):

Now, if you want me to derive that exponential function, you gotta ask nicely. It requires differential equations.

OpenStudy (anonymous):

I mean if you want me to derive: \[ F(t) = C\cdot 2^{kt} \]

OpenStudy (abb0t):

show off :P

OpenStudy (anonymous):

Yeah, but at the end of the day, I still get an inbox full of "Huh?" and no medals.

OpenStudy (anonymous):

ok this is what i was getting but for some reason it is showing up as wrong?

OpenStudy (anonymous):

What are you getting for \(k\) and \(C\)?

OpenStudy (anonymous):

Is your ending formula: \[\large F(t) = 2000\cdot 2^{2t} \]

OpenStudy (anonymous):

2000(2)^t/2

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

You solved for \(k\) wrong.

OpenStudy (anonymous):

\[ 2^{k(1/2)}=2\implies k/2 = 1 \implies k=2 \]

OpenStudy (anonymous):

@nyland1 Why is this question still open...?